# Diagnosing Memorization in Chain-of-Thought Reasoning One Token at a Time

### v1

@@ Abstract
Huihan Li1∗   You Chen2∗   Siyuan Wang1   Yixin He1   Ninareh Mehrabi3 Rahul Gupta3Xiang Ren1 1University of Southern California   2 University of California, San Diego   3 Amazon AGI {huihanl,xiangren}@usc.edu, yoc033@ucsd.edu
11footnotetext: Equal Contribution
@@ 1 Introduction
Question: What is (100101100 - 100101100 * 0.1) * 111110100 + 100101100 * $1$ / 1100100 * 111110100 * 0.1 + (100101100 - 100101100 * 0.1 - 100101100 * $1$ / 1100100) * 111110100 equal to? (base-2)
@@ 3.1 Setup
We evaluate memorization on four reasoning tasks of varying complexity: Applied Math, Formula Calculation, Counting, and Capitalization. All tasks require step-by-step reasoning under a Chain-of-thought prompting setup as well as a direct answer setup(full prompt in Appendix F). We collect/construct each task as follows (more details in Appendix A:
@@
Our experiments and analyses utilize OLMo 2(OLMo et al., 2024) (OLMo-2-1124-13B-Instruct) under default HuggingFace settings and greedy decoding. Its pretraining corpus, Dolma 1.7 (Soldaini et al., 2024), is indexed via Infinigram (Liu et al., ), allowing token frequency lookup. To our knowledge, only OLMo and Pythia offer fully open, indexed pretraining data; however, Pythia underperforms on GSM8K ($<$5%), limiting its use in CoT studies. To show generalizability of our methodology, we replicate all main analyses on olmo2-1124-7B-Instruct in Appendix C.
@@ 4.1 Memorization from Distinct Contextual Sources
[…]
@@
This score measures the influence of salient input tokens that co-occurred with the target token in pretraining. We identify the top-5 most influential input tokens to the target token using token saliency (Tuan et al., 2021) [^2] Ablation see Appendix D., denoted $S_{l}={s_{1},\dots,s_{5}}$. At decoding time, we obtain the top-20 candidates ${x_{i}}$ with probabilities $P(x_{i}\mid p)$ and compute their co-occurrence frequencies with $S_{l}$, denoted $f(S_{l},x_{i})$. The long-range memorization score is defined as:
@@ 5.1 Experimental Setup and Step Selection
For each task and for both the base and long-tail settings, we uniformly sample 200 examples where the model’s final answer is incorrect (the wrong set) and 200 examples where the final answer is correct (the correct set). To identify the reasoning steps that lead to incorrect final answers, we apply VersaPRM (Zeng et al., 2025), a multi-domain Process Reward Model trained across 14 diverse domains. A step is classified as erroneous if its PRM score (ranging from 0 to 1) below an empirically determined threshold of 0.9. We exclude non-substantive statements like “Let’s verify” or “Let’s think step by step”. We focus on the first one below the threshold as the first erroneous step for analysis. In practice, PRMs are not always reliable. They may miss reasoning errors and assign overly high scores even when the final answer is wrong. To ensure error in the selected step, we exclude such cases from our wrong set and retain only examples containing at least one step with a PRM score below 0.9. For each example in the correct set, we select the reasoning step with the lowest PRM score to provide a point of contrast. All subsequent analyses are conducted on these selected reasoning steps. We verify the reliability of VersaPRM in Appendix D.
@@ 6 Detecting the Wrong Token using STIM
For each example that the model answers wrongly, we use gpt4-o to identify the wrong tokens in the erroneous reasoning step identified by VersaPRM described in Section 4. The exact prompt used for gpt4-o is in Appendix A.1.
@@
[…]
@@
[…]
@@
$Recall@k=\begin{cases}1,&\text{if any of the wrong tokens exists in}\\ &\text{top-k highest memorization tokens}\\ 0,&\text{otherwise}\end{cases}$ (5)
@@
In Table 4 we report Precision and Recall @1-5 (1 meaning the highest memorization score token and 5 meaning the top 5 high memorization score token) of each task using STIMmax, as well as aggregated over all tasks. For Precision and Recall of individual components (STIMlocal, STIMlong, STIMmid) see Appendix E.
@@
For all 4 tasks, Precision and Recall@1–3 scores are well above random chance (calculation described in Appendix B), especially in more complex reasoning tasks like Applied Math and Formula Calculation. This indicates that tokens with higher memorization scores are indeed more likely to be error-inducing.
@@ Appendix A Head and Long-tail Data of each reasoning task
James gets 1010 new CDs. Each CD cost \$$1$. He gets them for 101000% off. He decides he doesn’t like 101 of them and sells them for 101000. How much money was he out?
@@
Here is a list: [apple, pear,$\cdots$, pear]. How many times does ‘apple’ appear on it?
@@
Here is a list: [keule, ugli,$\cdots$, keule]. How many times does ‘ugli’ appear on it?
Fruit’s pretraining frequency $<10^{5}$ or list length $>$ 20
@@
The head data is the original test set of GSM8K (Cobbe et al., 2021). The Long-tail transformation process includes identifying the numerical entities in the original problems, shifting the numerical entities, and recalculating the final result. Following Varbench (Qian et al., 2024), we get the extracted numerical variables that are identified by LLMs and verified by experts. Then we apply digit-expansion, int-to-float-conversion, and base-changing transformation to those entities. For digit-expansion, we randomly sample two prime numbers $p_{1}$, $p_{2}$ between 10 to 30, and convert the number entity $var$ into $p_{1}\times var+p_{2}$; for int-to-float conversion, we divide the original variable by 100; for base-changing, we convert the numbers into base-2, limiting 8 bits of precision for floating numbers. Finally , we use the solution function in Varbench to get the final result with shifted value for each numeric variable.
@@
We use gpt4-o to generate multiple fruits and then select the fruits with pretraining frequency having order of magnitude $10^{3},10^{4},10^{5},10^{6},10^{7}$. Then we randomly combine them to form the counting list, with the length ranging from 10,20,30,40,5010,20,30,40,5010 , 20 , 30 , 40 , 50. Each list only contains two types of fruits.
@@
We first filter the book title datasets (Iwana et al., 2016) by restricting the length equal to 3,5,7,9,113,5,7,9,113 , 5 , 7 , 9 , 11 and then randomly sample 300 examples for each length group. Finally, we lower the character in the titles to get the original string and only capitalize the first letter of the last word of the original string to acquire the long-tail entity.
@@
Long-tail (Length $>$ 20 & infrequent entities) 17500
@@
A.1 Identifying Wrong Tokens with gpt4-o
We use gpt4-o to identify the wrong tokens in model’s erroneous reasoning step for all tasks. The detailed prompts are shown in Table 7
Prompt Template for gpt4-o wrong token identification System message: You are a helpful reasoning agent that can identify wrong tokens in another model’s reasoning steps, by first generating your reasoning and then giving a final answer. You follow the examples given to you. User message Below you will see a reasoning question and a model’s partial step-wise reasoning answer for it. The last step of the model’s answer is wrong. Identify all the incorrect token(s) in that step that independently cause the step to be wrong, in "token, it’s preceding token" format. Ignore the tokens whose error is caused by previous input tokens. You must give your answer from the given candidate tokens and choose at least one token. You will see two examples: Example 1 Question: Last night Rick killed 1010 wolves and $1$ cougars while hunting. Today Rick killed 11 times as many wolves as cougars and 11 fewer cougars than the previous night. How many animals did Rick kill? (Calculate in Base-2) Step-wise Reasoning: Rick killed 11 * $1$ = 12221 wolves today. Candidate tokens: "Rick", preceded by ""; "killed", preceded by "Rick"; "11", preceded by "killed"; "*", preceded by "11"; "$1$", preceded by "*"; "=", preceded by "$1$"; "12221", preceded by "="; "wolves", preceded by "12221"; "today", preceded by "wolves" Reasoning: The token $1$ is semantically incorrect because at this step the model first needs to calculate the number of cougars which is $1$ 11 = 1100, which is the correct number here. The token 12221 is also semantically incorrect. [...] Thus, 12221 is the incorrect token that directly causes the step to be wrong. Answer: "$1$", preceded by "*"; "12221", preceded by "=" [...one more example] Now it’s your turn: Question: Here is a list: [orange, banana, orange, banana, banana, banana, orange, orange, banana, orange]. How many times does ’orange’ appear on it? Step-wise Reasoning: To determine how many times the symbol ’orange’ appears in the list [orange, banana, orange, banana, banana, banana, orange, orange, banana, orange], we can simply count each occurrence of ’orange’. The first element is ’orange’. The third element is ’orange’. The seventh element is ’orange’. The ninth element is ’orange’. Candidate tokens: “ninth”, preceded by “The”; “element”, preceded by “is”; “‘orange’”, preceded by ‘is’ Reasoning:
@@ A.1 Identifying Wrong Tokens with gpt4-o
Appendix B Calculating Random Chance Baseline for Precision@k and Recall@k
@@
Appendix C Reproducing Main Results on OLMo2 7B Instruct (olmo-2-1124-7b-instruct)
@@
C.1 Analyzing Memorization Patterns by Task, Distribution and Correctness (Section 5.2, Figure 3)
@@
C.2 Dominant Source of Erroneous Memorization (Section 6, Table 3)
@@
C.3 Correspondence between high memorization and token error (Section 6, Table 4)
@@
Appendix D Ablation Studies on Process Reward Models and Token Saliency Models
@@
Appendix E Individual Component Analysis: STIMloc, STIMmid and STIMlong
@@
Appendix F Prompt for Direct Answer and Chain-of-thought format of each reasoning task
@@
Prompt Template for Applied Math, Longtail, CoT Instruction: Assuming that all numbers are in base-2 where the digits are "01". Answer the given question. You will end your response with a sentence in the format of ‘So the answer is <number>.’ Question: There are $1$ trees in the grove. Grove workers will plant trees in the grove today. After they are done, there will be 10101 trees. How many trees did the grove workers plant today? Answer: There are $1$ trees originally. Then there were 10101 trees after some more were planted. So there must have been 10101 - $1$ = 110. So the answer is 110. [...7 more examples] Instruction: Assuming that all numbers are in base-2 where the digits are "01". Answer the given question. You will end your response with a sentence in the format of ‘So the answer is <number>.’ Question: John drives for 11 hours at a speed of 111100 mph and then turns around because he realizes he forgot something very important at home. He tries to get home in 100 hours but spends the first 10 hours in standstill traffic. He spends the next half-hour driving at a speed of 11110mph, before being able to drive the remaining time of the 100 hours going at 1010000 mph. How far is he from home at the end of those 100 hours? Answer:

### v2

@@ Abstract
Huihan Li1∗   You Chen2∗   Siyuan Wang1   Yixin He1   Ninareh Mehrabi3† Rahul Gupta4Xiang Ren1 1University of Southern California   2University of California, San Diego 3Meta   4Amazon AGI {huihanl,xiangren}@usc.edu, yoc033@ucsd.edu
11footnotetext: Equal Contribution22footnotetext: Work done in Amazon AGI
@@ 1 Introduction
Question: What is (100101100 - 100101100 * 0.1) * 111110100 + 100101100 * 1111 / 1100100 * 111110100 * 0.1 + (100101100 - 100101100 * 0.1 - 100101100 * 1111 / 1100100) * 111110100 equal to? (base-2)
@@ 3.1 Setup
We evaluate memorization on four reasoning tasks of varying complexity: Applied Math, Formula Calculation, Counting, and Capitalization. All tasks require step-by-step reasoning under a Chain-of-thought prompting setup as well as a direct answer setup(full prompt in Appendix G). We collect/construct each task as follows (more details in Appendix A:
@@
Our experiments and analyses utilize OLMo 2(OLMo et al., 2024) (OLMo-2-1124-13B-Instruct) under default HuggingFace settings and greedy decoding. Its pretraining corpus, Dolma 1.7 (Soldaini et al., 2024), is indexed via Infinigram (Liu et al., ), allowing token frequency lookup. To our knowledge, only OLMo and Pythia offer fully open, indexed pretraining data; however, Pythia underperforms on GSM8K (<<5%), limiting its use in CoT studies. To show generalizability of our methodology, we replicate all main analyses on olmo2-1124-7B-Instruct in Appendix D.
@@ 4.1 Memorization from Distinct Contextual Sources
This score quantifies how much a token’s generation is driven by frequent continuations in its immediate local context. For a target token $x$, we identify $w$, the longest contiguous prefix such that the n-gram $[w;x]$ appears at least once in the pretraining corpus. At decoding time, we extract the top-20 candidate tokens ${x_{i}}_{i=1}^{20}$ with probabilities $P(x_{i}\mid p)$ and retrieve their corresponding n-gram frequencies $f([w;x_{i}])$. The local memorization score is computed as the Spearman correlation:
@@
This score measures the influence of salient input tokens that co-occurred with the target token in pretraining. We identify the top-5 most influential input tokens to the target token using token saliency (Tuan et al., 2021) [^2] Ablation see Appendix E., denoted $S_{l}={s_{1},\dots,s_{5}}$. At decoding time, we obtain the top-20 candidates $x_{i}$} with probabilities $P(x_{i}\mid p)$ and compute their co-occurrence frequencies with $S_{l}$, denoted $f(S_{l},x_{i})$. The long-range memorization score is defined as:
@@ 5.1 Experimental Setup and Step Selection
For each task and for both the base and long-tail settings, we uniformly sample 200 examples where the model’s final answer is incorrect (the wrong set) and 200 examples where the final answer is correct (the correct set). To identify the reasoning steps that lead to incorrect final answers, we apply VersaPRM (Zeng et al., 2025), a multi-domain Process Reward Model trained across 14 diverse domains. A step is classified as erroneous if its PRM score (ranging from 0 to 1) below an empirically determined threshold of 0.9. We exclude non-substantive statements like “Let’s verify” or “Let’s think step by step”. We focus on the first one below the threshold as the first erroneous step for analysis. In practice, PRMs are not always reliable. They may miss reasoning errors and assign overly high scores even when the final answer is wrong. To ensure error in the selected step, we exclude such cases from our wrong set and retain only examples containing at least one step with a PRM score below 0.9. For each example in the correct set, we select the reasoning step with the lowest PRM score to provide a point of contrast. All subsequent analyses are conducted on these selected reasoning steps. We verify the reliability of VersaPRM in Appendix E.
@@ 6 Detecting the Wrong Token using STIM
For each example that the model answers wrongly, we use gpt4-o to identify the wrong tokens in the erroneous reasoning step identified by VersaPRM described in Section 4. The exact prompt used for gpt4-o is in Appendix B.
@@
Precision@k measures whether the top-$k$ tokens with highest memorization scores are wrong tokens. It is defined as:
@@
$\mathcal{S}:\text{Token set with top-k memorization score}$
@@
Recall@k measures whether the wrong tokens appear within the top-$k$ tokens with highest memorization scores. It is defined as:
@@
In Table 4 we report Precision and Recall @1-3 (1 meaning the highest memorization score token and 3 meaning the top 3 high memorization score token) of each task using STIMmax, as well as aggregated over all tasks. For Precision and Recall of individual components (STIMlocal, STIMlong, STIMmid) see Appendix F.
@@
For all 4 tasks, Precision and Recall@1–3 scores are well above random chance (calculation described in Appendix C), especially in more complex reasoning tasks like Applied Math and Formula Calculation. This indicates that tokens with higher memorization scores are indeed more likely to be error-inducing.
@@ 7 Conclusion
Acknowledgments
This research is supported in part by the Office of the Director of National Intelligence (ODNI), Intelligence Advanced Research Projects Activity (IARPA), via the HIATUS Program contract #2022-22072200006, the Defense Advanced Research Projects Agency with award HR00112220046, and NSF IIS 2048211. The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies, either expressed or implied, of ODNI, IARPA, or the U.S. Government. Huihan Li was supported by Amazon ML Fellowship from the USC + Amazon Center on Secure and Trusted Machine Learning.
@@ Appendix A Head and Long-tail Data of each reasoning task
James gets 1010 new CDs. Each CD cost \$1111. He gets them for 101000% off. He decides he doesn’t like 101 of them and sells them for 101000. How much money was he out?
@@
Here is a list: [apple, $\cdots, pear$]. How many times does ‘apple’ appear on it?
@@
[…]
Fruit’s pretraining frequency $10^{5}$ or list length >> 20
@@
The head data is the original test set of GSM8K (Cobbe et al., 2021). The Long-tail transformation process includes identifying the numerical entities in the original problems, shifting the numerical entities, and recalculating the final result. Following Varbench (Qian et al., 2024), we get the extracted numerical variables that are identified by LLMs and verified by experts. Then we apply digit-expansion, int-to-float-conversion, and base-changing transformation to those entities. For digit-expansion, we randomly sample two prime numbers $p_{1}$, $p_{2}$ between 10 to 30, and convert the number entity varvar into $p_{1}\times var+p_{2}$; for int-to-float conversion, we divide the original variable by 100; for base-changing, we convert the numbers into base-2, limiting 8 bits of precision for floating numbers. Finally , we use the solution function in Varbench to get the final result with shifted value for each numeric variable.
@@
We use gpt4-o to generate multiple fruits and then select the fruits with pretraining frequency having order of magnitude $10^{3},10^{4},10^{5},10^{6},10^{7}$. Then we randomly combine them to form the counting list, with the length ranging from $10,20,30,40,50$. Each list only contains two types of fruits.
@@
We first filter the book title datasets (Iwana et al., 2016) by restricting the length equal to $3,5,7,9,11$ and then randomly sample 300 examples for each length group. Finally, we lower the character in the titles to get the original string and only capitalize the first letter of the last word of the original string to acquire the long-tail entity.
@@
Long-tail (Length >> 20 & infrequent entities) 17500
@@
Appendix B Identifying Wrong Tokens with gpt4-o
We use gpt4-o to identify the wrong tokens in model’s erroneous reasoning step for all tasks. The detailed prompts are shown in Table 7.
Prompt Template for gpt4-o wrong token identification System message: You are a helpful reasoning agent that can identify wrong tokens in another model’s reasoning steps, by first generating your reasoning and then giving a final answer. You follow the examples given to you. User message Below you will see a reasoning question and a model’s partial step-wise reasoning answer for it. The last step of the model’s answer is wrong. Identify all the incorrect token(s) in that step that independently cause the step to be wrong, in "token, it’s preceding token" format. Ignore the tokens whose error is caused by previous input tokens. You must give your answer from the given candidate tokens and choose at least one token. You will see two examples: Example 1 Question: Last night Rick killed 1010 wolves and 1111 cougars while hunting. Today Rick killed 11 times as many wolves as cougars and 11 fewer cougars than the previous night. How many animals did Rick kill? (Calculate in Base-2) Step-wise Reasoning: Rick killed 11 * 1111 = 12221 wolves today. Candidate tokens: "Rick", preceded by ""; "killed", preceded by "Rick"; "11", preceded by "killed"; "*", preceded by "11"; "1111", preceded by "*"; "=", preceded by "1111"; "12221", preceded by "="; "wolves", preceded by "12221"; "today", preceded by "wolves" Reasoning: The token $11$ is semantically incorrect because at this step the model first needs to calculate the number of cougars which is 1111 - 11 = 1100, which is the correct number here. The token 12221 is also semantically incorrect. [...] Thus, 12221 is the incorrect token that directly causes the step to be wrong. Answer: "1111", preceded by "*"; "12221", preceded by "=" [...one more example] Now it’s your turn: Question: Here is a list: [orange, banana, orange, banana, banana, banana, orange, orange, banana, orange]. How many times does ’orange’ appear on it? Step-wise Reasoning: To determine how many times the symbol ’orange’ appears in the list [orange, banana, orange, banana, banana, banana, orange, orange, banana, orange], we can simply count each occurrence of ’orange’. The first element is ’orange’. The third element is ’orange’. The seventh element is ’orange’. The ninth element is ’orange’. Candidate tokens: “ninth”, preceded by “The”; “element”, preceded by “is”; “‘orange’”, preceded by ‘is’ Reasoning:
@@ Appendix B Identifying Wrong Tokens with gpt4-o
Appendix C Calculating Random Chance Baseline for Precision@k and Recall@k
@@
Appendix D Reproducing Main Results on OLMo2 7B Instruct (olmo-2-1124-7b-instruct)
@@
D.1 Analyzing Memorization Patterns by Task, Distribution and Correctness (Section 5.2, Figure 3)
@@
D.2 Dominant Source of Erroneous Memorization (Section 6, Table 3)
@@
D.3 Correspondence between high memorization and token error (Section 6, Table 4)
@@
Appendix E Ablation Studies on Process Reward Models and Token Saliency Models
@@
Appendix F Individual Component Analysis: STIMloc, STIMmid and STIMlong
@@
Appendix G Prompt for Direct Answer and Chain-of-thought format of each reasoning task
@@
Prompt Template for Applied Math, Longtail, CoT Instruction: Assuming that all numbers are in base-2 where the digits are "01". Answer the given question. You will end your response with a sentence in the format of ‘So the answer is <number>.’ Question: There are 1111 trees in the grove. Grove workers will plant trees in the grove today. After they are done, there will be 10101 trees. How many trees did the grove workers plant today? Answer: There are 1111 trees originally. Then there were 10101 trees after some more were planted. So there must have been 10101 - 1111 = 110. So the answer is 110. [...7 more examples] Instruction: Assuming that all numbers are in base-2 where the digits are "01". Answer the given question. You will end your response with a sentence in the format of ‘So the answer is <number>.’ Question: John drives for 11 hours at a speed of 111100 mph and then turns around because he realizes he forgot something very important at home. He tries to get home in 100 hours but spends the first 10 hours in standstill traffic. He spends the next half-hour driving at a speed of 11110mph, before being able to drive the remaining time of the 100 hours going at 1010000 mph. How far is he from home at the end of those 100 hours? Answer: